OpenStudy (anonymous):

solve the equation 4y-8-2y+5=0

OpenStudy (ghazi):

4y-2y-8+5=0 2y-3=0 2y=3 y=3/2

OpenStudy (anonymous):

thank you soo much

OpenStudy (ghazi):

:) you're welcome